Solve using the quadratic formula:

x2 + 4x + 3 = 0


1
Expert's answer
2021-02-22T18:31:40-0500

For the quadratic equation

"ax^2+bc+c=0"

the formula for the roots is


"x=\frac{-b± \sqrt{b^2-4ac}}{2a}"


For. "x^2+4x+3=0"


"a=1, b=4, c= 3"


Hence


"x=\frac{-4± \sqrt{4^2-(4×1×3)}}{2×1}"


"x=\frac{-4± \sqrt{16-12}}{2}"


"x=\frac{-4± \sqrt{4}}{2}"


"x=\frac{-4±2}{2}"


"x"1"=-1"

"x" 2"=-3"


Hence

"x" is either -1 or -3